# Pareto-Optimal Learning from Preferences with Hidden Context

### v1

@@ C.1 Synthetic Experiments
We follow the experimental procedure of Siththaranjan et al. [50] in generating preferences, except modify their code such that we ensure that annotator identity is held constant for each preference. We use last layer finetuning on a neural network that is randomly initialized. We did not include any pre-training here to ensure that we are not pushing our reward models towards any modes before starting to train. We use a batch size of 2048 preferences, a step size of 0.1 and 10000 steps of MCMC for B-Rex. For POPL, we use a population size of 100 and a generation count of 100. We use a $\beta$ (confidence) value of 10, although have found that changing this value does not significantly affect B-REx’s performance.
C.2 Gridworld Model Experiments
@@
For the Gridworld model experiments, we base our environment on the Minigrid package [16]. Demonstrations were generated by rolling out many checkpointed policies at different levels of performance, trained using Proximal Policy Optimization (PPO). Then, these demonstrations were annotated based on a high performing policy’s action selection probabilities.
@@
For MultiCPL and POPL, we use behavior cloning directly on the demonstrations for 1000 iterations with a batch size of 64 and a learning rate of 0.001 with the Adam optimizer [36] as pretraining. The model architecture was a simple convolutional neural network that takes input from the agent’s view window, and has a single fully connected layer with 128 nodes to output the 7 actions from the environment. For both MultiCPL and POPL, we use last layer fine-tuning. For MultiCPL, we use the CPL objective, a learning rate of 0.001, where each model in a population of 500 models is trained for 20 iterations. For POPL, we use a learning rate of 0.2, and 1000 total steps. We sample 640 preferences every 10 iterations (as we can cache the last layer features for this examples for improved performance), and sub-sample a batch of size 64 for each step of lexicase selection. For a fair comparison between these two approaches, we approximately hold constant total wall clock time on the same hardware. Given a final population of policies generated by POPL or MultiCPL, we select the top 10 models for each hidden context class as the catered policy for that group.
C.3 Language Model Experiments
@@
In the LLM experiments, we assess the performance of reward learning by examining preference accuracy on the test set. To investigate vulnerabilities to jailbreak, we analyze pairs of responses to jailbreak prompts designed by Wei et al. [53] to deceive the model into giving a harmful response. We calculate the percentage of prompts where it assigns a higher reward to the jailbroken response (”jailbreak rate”). Additionally, we evaluate the reward function’s ability to assess helpfulness on non-harmful prompts, i.e., the reward function predicts higher rewards on the more helpful response. We compare our method to normal RLHF with an LLM-based preference model, Bayesian Reward Extrapolation (B-REx), and distributional preference learning (DPL). DPL methods predict parameters of the distribution of reward values for each response, rather than a single reward value, in order to better account for hidden context in human preferences.
@@
For standard RLHF, we use the pre-trained Llama-2-7b [52] preference model by Siththaranjan et al. [50], which is fine-tuned on the HH-RLHF dataset using LoRA [33]. We implement B-REx by performing linear reward extrapolation on the last layer of the pre-trained Llama-2-7b preference model. Following the B-REx implementation in [11], we run 200,000 steps of MCMC with a step size of 0.05. We use a burn-in of 5000 and a skip every 20 samples to reduce auto-correlation. For POPL, we run lexicase selection for 100 generations with a population size of 1000, and randomly sample 100 reward functions in the last generation.

### v2

@@ C.1 Synthetic Experiments
We follow the experimental procedure of Siththaranjan et al. (2023) in generating preferences, except modify their code such that we ensure that annotator identity is held constant for each preference. We use last layer finetuning on a neural network that is randomly initialized. We did not include any pre-training here to ensure that we are not pushing our reward models towards any modes before starting to train. We use a batch size of 2048 preferences, a step size of 0.1 and 10000 steps of MCMC for B-Rex. For POPL, we use a population size of 100 and a generation count of 100. We use a $\beta$ (confidence) value of 10, although have found that changing this value does not significantly affect B-REx’s performance.
C.2 Minigrid Experiments
@@
For the Gridworld model experiments, we base our environment on the Minigrid package Chevalier-Boisvert et al. (2023). Demonstrations were generated by rolling out many checkpointed policies at different levels of performance, trained using Proximal Policy Optimization (PPO). Then, these demonstrations were annotated based on a high performing policy’s action selection probabilities.
@@
For MultiCPL and POPL, we use behavior cloning directly on the demonstrations for 1000 iterations with a batch size of 64 and a learning rate of 0.001 with the Adam optimizer Kingma & Ba (2014) as pretraining. The model architecture was a simple convolutional neural network that takes input from the agent’s view window, and has a single fully connected layer with 128 nodes to output the 7 actions from the environment. For both MultiCPL and POPL, we use last layer fine-tuning. For MultiCPL, we use the CPL objective, a learning rate of 0.001, where each model in a population of 500 models is trained for 20 iterations. For POPL, we use a learning rate of 0.2, and 1000 total steps. We sample 640 preferences every 10 iterations (as we can cache the last layer features for this examples for improved performance), and sub-sample a batch of size 64 for each step of lexicase selection. For a fair comparison between these two approaches, we approximately hold constant total wall clock time on the same hardware. Given a final population of policies generated by POPL or MultiCPL, we select the top 10 models for each hidden context class as the catered policy for that group.
C.3 Metaworld experiments
For the Metaworld robotics benchmark (Yu et al., 2019), we create augmented reward functions with greater emphasis on speed or safety, respectively. For the speed reward function, we add a penalty of $\frac{10}{T}$, where $T$ is the maximum timesteps allowed for that environment, for every timestep until the goal (as defined by the metaworld environment) is met. For the safety reward function, we add a penalty of $10\cdot\left\|\Omega\right\|_{2}$ where $\Omega$ is the angular velocity of all the robot’s joints. We also include, for each group, the reward from the other group, weighted with $0.1$ instead of $10$.
We generated demonstrations by training optimal policies on each task studied using Proximal Policy Optimization (Schulman et al., 2017) with the Stable Baselines package (Raffin et al., 2021). Every 100,000 steps, we cached the policy parameters to be used to generate sub-optimal performance. We train one policy on each reward function for a total of 1 million timesteps. We then roll out the policies at each checkpoint to generate $600$ demonstrations, that are used to select snippets of length $150$ that are ranked using log-likelihoods of the trajectory snippets under the optimal policy. These preferences are fed to the preference learning system.
The experiments follow a very similar outline to the Minigrid experiments outlined in Appendix C.2 above. All frameworks use the same network architecture: A simple two layer Neural Network with 1024 hidden nodes. For the button-press-v2 env, for example, this policy has 35 input nodes, 1024 hidden nodes, and 4 output nodes, with ReLU activation at the hidden layer. For both MultiCPL and POPL, we pre-train with behavior cloning directly from the demonstrations for 1000 iterations at a batch size of 16 and learning rate of $0.001$. We use last layer finetuning for both POPL and MultiCPL. For MultiCPL, we use the CPL objective, and train the last layer using a batch size of $16$, learning rate of $0.001$, and for $50$ iterations each. For POPL, we sample $512$ preferences every 25 iterations, and sub sample a batch of size $256$ to use for lexicase selections. We use a Gaussian mutation with mean 00 and standard deviation of $0.01$ to mutate our policies at each step.
C.4 Language Model Experiments
@@
In the LLM experiments, we assess the performance of reward learning by examining preference accuracy on the test set. To investigate vulnerabilities to jailbreak, we analyze pairs of responses to jailbreak prompts designed by Wei et al. (2024) to deceive the model into giving a harmful response. We calculate the percentage of prompts where it assigns a higher reward to the jailbroken response (“jailbreak rate”). Additionally, we evaluate the reward function’s ability to assess helpfulness on non-harmful prompts, i.e., the reward function predicts higher rewards on the more helpful response. We compare our method to normal RLHF with an LLM-based preference model, Bayesian Reward Extrapolation (B-REx), and distributional preference learning (DPL). DPL methods predict parameters of the distribution of reward values for each response, rather than a single reward value, in order to better account for hidden context in human preferences.
@@
For standard RLHF, we use the pre-trained Llama-2-7b (Touvron et al., 2023) preference model by Siththaranjan et al. (2023), which is fine-tuned on the HH-RLHF dataset using LoRA (Hu et al., 2022). We implement B-REx by performing linear reward extrapolation on the last layer of the pre-trained Llama-2-7b preference model. Following the B-REx implementation in (Brown et al., 2020a), we run 200,000 steps of MCMC with a step size of 0.05. We use a burn-in of 5000 and a skip every 20 samples to reduce auto-correlation. For POPL, we run lexicase selection for 100 generations with a population size of 1000, and randomly sample 100 reward functions in the last generation.